A population consists of five numbers 1, 2, 3, 4, 5 and 6. Suppose samples of size 2 are drawn from this population.


(a) Find the mean and variance of the population


(b) Describe the sampling distribution of the sample means.


(c) Find the mean and variance of the sampling distribution of the sample means.



1
Expert's answer
2022-05-27T11:41:58-0400

1.

 We have population values "1,2,3,4,5,6" population size "N=6"

"\mu=\dfrac{1+2+3+4+5+6}{6}=3.5"






"\sigma^2=\dfrac{1}{6}((1-3.5)^2+(2-3.5)^2+(3-3.5)^2""+(4-3.5)^2+(5-3.5)^2+(6-3.5)^2)=2.92"



The number of possible samples which can be drawn without replacement is





"\dbinom{N}{n}=\dbinom{6}{2}=15"






"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 1,2 & 1.5 \\
 \hdashline
 2 & 1,3 & 2 \\
 \hdashline
 3 & 1,4 & 2.5\\
 \hdashline
 4 & 1,5 & 3 \\
 \hdashline
 5 & 1,6 & 3.5 \\
 \hdashline
 6 & 2,3 & 2.5 \\
 \hdashline
 7 & 2,4 & 3 \\
 \hdashline
 8 & 2,5& 3.5 \\
 \hdashline
 9 & 2,6 & 4 \\
 \hdashline
 10 & 3,4 &3.5\\
\hdashline
 11 & 3,5 & 4 \\
 \hdashline
 12 & 3,6 & 4.5 \\
 \hdashline
 13&4,5 & 4.5 \\
 \hdashline
 14 & 4,6 & 5 \\
 \hdashline
 15 & 5,6 &5.5 \\
 \hline
\end{array}"



2.

The sampling distribution of the sample means.




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
& \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 & 1.5 & 1 & 1/15 &0.1& 0.15 \\
 \hdashline
 & 2& 1 & 1/15 & 0.13 & 0.266 \\
 \hdashline
 & 2.5& 2 & 2/15 & 0.333& 0.8333 \\
 \hdashline
 & 3 & 2 & 2/15 & 0.4& 1.2 \\
 \hdashline
 & 3.5 & 3 & 3/15 & 0.7 & 2.45 \\
 \hdashline
 & 4 & 2 & 2/15& 0.533 & 2.133 \\
 \hdashline
 & 4.5 & 2 & 2/15 & 0.6 & 2.7 \\
 \hdashline
 & 5 & 1 & 1/15 & 0.333 & 1.666 \\
 \hdashline
 & 5.5 & 1 & 1/15 & 0.366 & 2.0166\\
 \hdashline
 
 Total & & 15 & 1 & 3.5 & 13.415\\ \hline
\end{array}"




4.



"E(\bar{X})=\sum\bar{X}f(\bar{X})=3.5"


The mean of the sampling distribution of the sample means is equal to the mean of the population.




"E(\bar{X})=\mu_{\bar{X}}=3.5=\mu"




"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"






"={13.415}-(3.5)^2=1.165"